# The PLATO field selection process

### v1

@@ II. Characterization of LOPS2, the first long-pointing field
PLAnetary Transits and Oscillations of stars (PLATO) is an ESA M-class mission to be launched by the end of 2026 to discover and characterize transiting planets around bright and nearby stars, and in particular habitable rocky planets hosted by solar-like stars. Over the mission lifetime, an average of 8% of the science data rate will be allocated to Guest Observer programs (GOs) selected by ESA through public calls, hence it is essential for the community to know in advance where the observing fields will be located. In a previous paper, we identified two preliminary long-pointing fields (LOPN1 and LOPS1) for PLATO, respectively in the northern and southern hemisphere. Here we present LOPS2, a slightly adjusted version of the southern field that has recently been selected by the PLATO Science Working Team as the first field to be observed by PLATO for at least two continuous years, following the scientific requirements. In this paper, we describe the astrophysical content of LOPS2 in detail, including known planetary systems, bright/variable/binary stars, clusters and synergies with other current and future facilities.
@@ 1 Introduction
PLATO (PLAnetary Transits and Oscillations of stars; Rauer et al. 2024) is an ESA M-class mission primarily designed to detect a large number of planetary systems hosted by nearby and bright stars through the transit technique. Among many other aims, the goal of this mission is the discovery and confirmation of Earth-like planets, i. e., habitable rocky planets hosted by solar-type dwarfs (Heller et al., 2022; Matuszewski et al., 2023). Very accurate stellar parameters, including stellar age at 10% accuracy [^1] For solar-like stars., will be extracted from the light curves themselves through asteroseismological analysis (Cunha et al., 2021; Bétrisey et al., 2023). This will enable us to put the discovered planetary systems in a consistent evolutionary context. Furthermore, the follow-up and confirmation of candidate planets discovered by PLATO is part of the mission through the Ground-based Observation Program (GOP) of the PLATO Consortium. We refer the reader to Rauer et al. (2024) for a recent, detailed review of the mission.
@@
During its nominal 4-year mission, PLATO will continuously monitor one or two pre-selected fields for at least two years each (during the so-called LOP: Long-duration Observing Phase). The mission has been designed to have the capability to observe additional fields for shorter intervals, at least two months each (SOP; Short-duration Observing Phase). The 24 “normal” cameras of PLATO (NCAMs) are not all co-aligned, but rather split in four groups with an angular offset (9.2) with respect to the satellite bore-sight (Pertenais et al., 2021). As a consequence, the overall field of view (FOV) of about 2 14921492\,1492 149 $\textrm{deg}^{2}$ is covered by a variable number of NCAMs, from six to 24, depending on the specific line of sight considered, as illustrated by Fig. 1. In terms of sky area, approximately 325 deg2 will be covered by 24 NCAMs, 153 deg2 by 18 NCAMs, 847 deg2 by 12 NCAMs and 824 deg2 by 6 NCAMs (assuming a corrected FOV of $19.2^{\circ}$ and CCD gaps of 1.3 mm; Pertenais et al. 2021) although the exact numbers could slightly change according to the actual optical performances once in flight.
@@
Due to telemetry constraints, we will not be able to download full-frame images at high cadence [^2] Full-frame images will be acquired during the commissioning phase, at least once every 90 days.. Instead, PLATO will perform a significant part in the photometric analysis onboard. Only light curves, centroid data and a number of stamp-like “imagettes” will be transmitted back to Earth, meaning that targets have to be selected in advance. PLATO targets are divided into four main samples, summarized in Table 1, with different requirements on magnitude, spectral type and noise-to-signal ratio (NSR): three of such samples are composed by main-sequence or subgiant stars with spectral type from F5 to K7 (P1, P2, P5), while a fourth one contains only M dwarfs (P4). The PLATO Scientific Requirement Document [^3] ESA reference: PTO-EST-SCI-RS-0150. (SciRD) identifies P1 ($V<11$, $\textrm{NSR}<50$ ppm in one hour) as the highest-priority sample, and sets a requirement of at least 15 0001500015\,00015 000 P1 targets to be observed during the LOPs. We refer to the SciRD, Nascimbeni et al. (2022) and Montalto et al. (2021) (hereafter quoted as N22 and M21) for a complete summary of the definitions and requirements on the PLATO samples.
@@
Figure 1: Geometry of the PLATO field FOV. In this figure, the field is centered at the origin (0,0) of a generic spherical reference frame (units are degrees; projection is orthographic). The number of “normal” cameras covering a given line of sight is color-coded. The four blue shades, from dark to light, map the regions covering 325/153/847/824 deg2 observed respectively with 24, 18, 12, six cameras (and corresponding to four, three, two, and one group of six co-pointing telescopes each).
@@
PLATO is currently on schedule to be launched at the end of 2026, and to start its routine scientific operations in mid-2027. Most raw and calibrated data products from PLATO will be made publicly available as soon as possible [^4] With the exception of GO programs and reserved targets allocated to the Plato Mission Consortium, within their proprietary time.. The astronomical community has the opportunity to be deeply involved at different stages: preparation, data analysis, and follow-up. PLATO data products will be released quarterly. Each set of data from a three-month observation period will be made available approximately six month later. For around 20 0002000020\,00020 000 bright targets prioritized for ground-based follow-up observations, the data products will be delivered roughly 1.2 years after the corresponding three-month observation period. On top of that, over the mission lifetime, an average of 8% of the PLATO science data rate will be allocated to Guest Observer programs (GOs) selected by ESA through open calls, so it is essential to select and communicate each field that is going to be pointed at with a proper time margin.
@@
In a previous work (Paper I; N22), we described all the steps of the process that led to the identification of two provisional fields (labeled LOPN1 and LOPS1) as candidates for the LOP. Back then, it was still unknown which one would be selected as the first one. Also, it was anticipated that the final LOP fields could have had small changes in position and/or rotation angle. In this paper, we present LOPS2, the long-pointing field finally chosen by the ESA PLATO Science Working Team (PSWT) as the first field to be observed by PLATO. After summarizing the fine-tuning process and the general properties of LOPS2 in Section 2, we review and discuss its astrophysical content in Section 3. Finally, the conclusions and some prospects for the future work to be done are given in Section 4. A glossary of the most common acronyms used throughout this work is compiled in Table 3 in the Appendix.
@@ 2 The choice of the first long-pointing field
  • Rotation angle: is there any technical or scientific advantage to choose a different rotation angle?
  • Synergies: which field is the most interesting one for possible synergies with other facilities, in particular with regard to the radial-velocity follow-up process, considering the time frame of the mission?
In what follows we will discuss the impact of these three key ingredients on the final decision.
@@ 2.1 Duty cycle
  • the prioritization metric as defined in Paper I is 1% smaller (0.980 vs. 0.990), i.e., the difference is negligible;
@@ 2.2 Rotation angle
It is worth considering whether there is any other reason to rotate the field, for instance due to specific astrophysical objects we would like to include, or avoid, in the LOP field. As for the former, we verified that 1) the number of known planetary systems is not significantly affected by $\varphi$ for LOPN1, and is actually maximized at $\varphi=0^{\circ}$-$10^{\circ}$ for LOPS2 due to an overdensity of TESS-discovered planets lying in the Galactic south-west corner of the field (as we will show in Section 3.4); and 2) no open cluster among the ones of potential interest for PLATO (i. e., nearby, sparse, and mature) slips in or out of the field for any rotation angle choice, as it will be shown in Section 3 for LOPS2. More in general, the range of Galactic latitudes probed is always $0^{\circ}<|b|\lesssim 50^{\circ}$ regardless of $\varphi$, and therefore any property correlated with $b$, such as the amount of stellar crowding and the diversity of stellar populations probed (in terms of age, kinematics, metallicity, etc.) is approximately the same.
@@
Finally, it makes sense to investigate whether there is any way to rotate the field in order to move any extremely bright star ($V<1$) off the silicon. Paper I identified Deneb and Vega as the brightest sources on LOPN1, and Sirius and Canopus for LOPS1. Moving from LOPS1 to LOPS2 also implies that Sirius is now left outside the focal plane for any choice of $\varphi$, while there is no consequence on the other three stars, that are unavoidable and always monitored with the same number of NCAMs (18-24 for Vega, 6 for Deneb, 24 for Canopus) for almost any possible rotation angle [^6] It could be possible in principle to fit Deneb within a $\sim 30^{\prime}$ CCD gap of the northern field at $\varphi\simeq 11^{\circ}$, but the technical feasibility of such a choice and its possible unintended consequences (see for instance Porterfield et al. 2016) need to be assessed. This will be done at a further stage, should LOPN1 be scheduled for observation.. Fig. 15 (but also Section 3.2 below) investigates this point a bit further, looking at how the number of bright stars (up to $V<3$) in LOPS2 changes as a function of $\varphi$, at $10^{\circ}$ intervals from $0^{\circ}$ to $80^{\circ}$ (the symmetry of the PLATO field makes it invariant after $90^{\circ}$ rotations). The nominal $\varphi=0^{\circ}$ case is the only one for which all the five $21<V<$ 2 stars ($\delta$ CMa, $\epsilon$ CMa, $\gamma_{2}$ Vel, $\delta$ Vel, $\epsilon$ Car) always land on 6 NCAMs regions, while for instance four of them fall on 12 NCAMs in the $\varphi=40$-$50^{\circ}$ case. Finally, also the distribution of $32<V<$ 3 stars does not change significantly as a function of $\varphi$.
@@
In summary, we conclude that there is no reason to change the rotation angle of both LOPN1 and LOPS2 from its initial value of $\varphi=0^{\circ}$. We will implicitly assume zero rotation (i.e., one side of the field almost tangent to the Galactic plane) hereafter, as plotted in the all-sky projection in Fig. 2.
@@
Figure 2: All-sky Aitoff projection in Galactic coordinates of LOPS2 and LOPN1, showing the formal constraints for the selection of the PLATO LOP fields and the synergies with other missions. The two pink circles represent the $|\beta|>63^{\circ}$ technical requirement for the center of the LOP fields (“allowed region”), implying that the overall envelopes of every allowed field choice are two ecliptic caps at $|\beta|\gtrsim 38^{\circ}$ (green circles). LOPS2 (lower left) and LOPN1 (upper right) are plotted with blue shades according to the number of co-pointing cameras, as in Fig. 1. The footprints of CoRoT (red), Kepler (magenta), and K2 (green) are over-plotted together with the TESS continuous viewing zone at $|\beta|\gtrsim 78^{\circ}$ (yellow circle). The background gray layer is color-coded according to the areal density of $G<13.5$ stars from Gaia DR3. The celestial equator and poles are marked with a red line and crosses, respectively. This sky chart is also plotted in equatorial and Ecliptic coordinates in Fig. 17-18.
@@ 2.3 Synergies and follow-up
As for the previous space-based transit surveys, we first remind that no overlap between a LOP field and K2 or CoRoT (Howell et al., 2014; Auvergne et al., 2009) is possible at all, since the latter fields are all located close to the ecliptic by design, inaccessible during the LOP phase (Fig. 2). The TESS continuous viewing zone (CVZ) at $|\beta|>78^{\circ}$, on the other hand, is monitored at 100% in LOPN1 and $\gtrsim 90\%$ in LOPS2, mostly in six- and 12-camera regions. Finally, the Kepler field is fully included in the LOPN1, almost entirely with at least 12 NCAMs and up to 24. While the scientific case of following-up the Kepler TTV systems with PLATO is very solid (Jontof-Hutter et al., 2021a, b), we emphasize that it should not be seen as a main driver of the mission, but rather as a potential source of interesting additional science [^7] It is worth mentioning that the follow-up of the Kepler field is part of the science goals of the proposed Earth-2.0 mission (Ge et al., 2022).. Indeed, the last word is on the follow-up availability and timeliness, since the confirmation of the PLATO candidates and eventual mass determination is part of the mission products (and goals) through the Ground-based Observation Program (GOP) of the PLATO Consortium.
@@
It is worth mentioning two other important survey projects in the South that could have interesting implications for the PLATO preparation and/or follow-up: LSST (Ivezić et al., 2019) at the Vera C. Rubin Observatory, that will map the whole southern hemisphere (including LOPS2) beginning in 2025, and 4MOST at the VISTA telescope (de Jong et al., 2012), starting its consortium surveys also in 2024.
@@ 2.4 LOPS2 as first LOP field
 [^8] The rows give: coordinates for the field center in equatorial, Galactic and ecliptic reference frames, and number of targets available in the P1-P2-P4-P5 samples. The latter are calculated from PIC 2.0.0 assuming the “EOL 22” scenario (see Section 2.4).
Figure 3: LOPS2 footprint. Left panel: orthographic projection in Galactic coordinates centered on LOPS2. The field is color-coded according to the number of NCAMs as in Fig. 1. The yellow circle at $\beta=-78^{\circ}$ approximates the southern TESS continuous viewing zone (CVZ), while the magenta circle is centered on the Large Magellanic cloud. Right panel: as in left panel, but in equatorial coordinates.
@@
Figure 4: Interstellar extinction properties of LOPS2. Left panel: Average $V$-band extinction $A_{V}$ of the P1+P2 targets from PIC 2.0.0, calculated over HEALPix level-5 cells and represented in a color scale. Right panel: same, but for the P5 sample and over a finer level-6 grid. The high-extinction bump at $A_{V}\simeq 0.3$ is due to the closest parts of the Vela-Puppis region (Cantat-Gaudin et al., 2019).

### v3

@@ II. Characterization of LOPS2, the first long-pointing field
PLAnetary Transits and Oscillations of stars (PLATO) is an ESA M-class mission to be launched by the end of 2026 to discover and characterize transiting planets around bright and nearby stars, and in particular habitable rocky planets hosted by solar-like stars. Over the mission lifetime, an average of 8% of the science data rate will be allocated to Guest Observer programs selected by ESA through public calls. Hence, it is essential for the community to know in advance where the observing fields will be located. In a previous paper, we identified two preliminary long-pointing fields (LOPN1 and LOPS1) for PLATO, respectively in the northern and southern hemispheres. Here we present LOPS2, a slightly adjusted version of the southern field that has recently been selected by the PLATO Science Working Team as the first field to be observed by PLATO for at least two continuous years, following the scientific requirements. In this paper, we describe the astrophysical content of LOPS2 in detail, including known planetary systems, bright stars, variables, binary stars, star clusters, and synergies with other current and future facilities.
@@ 1 Introduction
PLAnetary Transits and Oscillations of stars (PLATO; Rauer et al. 2024) is an ESA M-class mission primarily designed to detect a large number of planetary systems hosted by nearby and bright stars through the transit technique. Among many other aims, the goal of this mission is the discovery and confirmation of Earth-like planets, that is, habitable rocky planets hosted by solar-type dwarfs (Heller et al., 2022; Matuszewski et al., 2023). Very accurate stellar parameters, including stellar age at 10% accuracy, [^1] For solar-like stars. will be extracted from the light curves themselves through asteroseismological analysis (Cunha et al., 2021; Bétrisey et al., 2023). This will enable the discovered planetary systems to be placed in a consistent evolutionary context. Furthermore, follow-up and confirmation of candidate planets discovered by PLATO is part of the mission through the Ground-based Observation Program (GOP) of the PLATO Consortium. We refer the reader to Rauer et al. (2024) for a recent, detailed review of the mission.
@@
During its nominal four-year mission, PLATO will continuously monitor one or two pre-selected fields for at least two years each during the so-called Long-duration Observation Phase (LOP). The mission has been designed to have the capability to observe additional fields for shorter intervals, at least two months each (Step-and-stare Observation Phase; SOP). The 24 “normal” cameras (NCAMs) of PLATO are not all co-aligned, but rather split into four groups with an angular offset (9.2) with respect to the satellite bore-sight (Pertenais et al., 2021). As a consequence, the overall field of view (FOV) of about 2 14921492\,1492 149 $\textrm{deg}^{2}$ is covered by a variable number of NCAMs, from six to 24, depending on the specific line of sight considered, as illustrated by Fig. 1. In terms of sky area, approximately 325 deg2 will be covered by 24 NCAMs, 153 deg2 by 18 NCAMs, 847 deg2 by 12 NCAMs, and 824 deg2 by six NCAMs (assuming a corrected FOV of $19.2^{\circ}$ and CCD gaps of 1.3 mm; Pertenais et al. 2021), although the exact numbers could slightly change according to the actual optical performances once in flight.
@@
Due to telemetry constraints, PLATO will not be able to download full-frame images at high cadence. [^2] Full-frame images will be acquired during the commissioning phase, at least once every 90 days. Instead, it will perform a significant part of the photometric analysis onboard. Only light curves, centroid data, and a number of stamp-like “imagettes” will be transmitted back to Earth, meaning that targets have to be selected in advance. PLATO targets are divided into four main samples, summarized in Table 1, with different requirements on magnitude, spectral type, and noise-to-signal ratio (NSR); three of such samples are composed by main-sequence or subgiant stars with a spectral type from F5 to K7 (P1, P2, P5), while a fourth one contains only M dwarfs (P4). The PLATO Scientific Requirement Document [^3] ESA reference: PTO-EST-SCI-RS-0150. (SciRD) identifies P1 ($V<11$, $\textrm{NSR}<50$ ppm in one hour) as the highest-priority sample, and sets a requirement of at least 15 0001500015\,00015 000 P1 targets to be observed during the LOPs. We refer to the SciRD, Nascimbeni et al. (2022), and Montalto et al. (2021, hereafter quoted as N22 and M21) for a complete summary of the definitions and requirements on the PLATO samples.
@@
Figure 1: Geometry of the PLATO field FOV. In this figure, the field is centered at the origin (0,0) of a generic spherical reference frame (units are degrees; projection is orthographic). The number of “normal” cameras covering a given line of sight is color coded. The four blue shades, from dark to light, map the regions covering 325/153/847/824 deg2 observed respectively with 24, 18, 12, six cameras (and corresponding to four, three, two, and one group of six co-pointing telescopes each).
@@
PLATO is currently on schedule to be launched at the end of 2026 and to start its routine scientific operations in mid-2027. Most raw and calibrated data products from PLATO will be made publicly available as soon as possible. [^4] With the exception of GO programs and reserved targets allocated to the Plato Mission Consortium, within their proprietary time. The astronomical community has the opportunity to be deeply involved at different stages: preparation, data analysis, and follow-up. PLATO data products will be released quarterly. Each set of data from a three-month observation period will be made available approximately six months later. For around 20 0002000020\,00020 000 bright targets prioritized for ground-based follow-up observations, the data products will be delivered roughly 1.2 years after the corresponding three-month observation period. On top of that, over the mission lifetime, an average of 8% of the PLATO science data rate will be allocated to Guest Observer (GO) programs selected by ESA through open calls, so it is essential to select and communicate each field that is going to be pointed at with a proper time margin.
@@
In a previous work (Paper I; N22), we described all the steps of the process that led to the identification of two provisional fields (labeled LOPN1 and LOPS1) as candidates for the LOP. Back then, it was still unknown which field would be selected as the first one. Also, it was anticipated that the final LOP fields could have had small changes in position and/or rotation angle. In this paper, we present LOPS2, the long-pointing field finally chosen by the ESA PLATO Science Working Team (PSWT) as the first field to be observed by PLATO. After summarizing the fine-tuning process and the general properties of LOPS2 in Section 2, we review and discuss its astrophysical content in Section 3. Finally, the conclusions and some prospects for the future work to be done are given in Section 4. A glossary of the most common acronyms used throughout this work is compiled in Table 3 in the Appendix.
@@ 2 The choice of the first long-pointing field
  • Rotation angle: the choice of a different rotation angle should be investigated for any technical or scientific advantage.
  • Synergies: the field with the most interesting synergies with other facilities should be given higher priority, in particular with regard to the radial-velocity follow-up process, considering the time frame of the mission.
In what follows, we discuss the impact of these three key ingredients on the final decision. The final decision is presented and discussed in Section 2.4.
@@ 2.1 Duty cycle
  • the prioritization metric as defined in Paper I is 1% smaller (0.980 vs. 0.990), that is, the difference is negligible;
@@ 2.2 Rotation angle
It is worth considering whether there is any other reason to rotate the field, for instance due to specific astrophysical objects we would like to include, or avoid, in the LOP field. As for the former, we verified that 1) the number of known planetary systems is not significantly affected by $\varphi$ for LOPN1, and is actually maximized at $\varphi=0^{\circ}$-$10^{\circ}$ for LOPS2 due to an overdensity of TESS-discovered planets lying in the galactic southwest corner of the field (as we show in Section 3.4); and 2) no open cluster among the ones of potential interest for PLATO (i. e., nearby, sparse, and mature) slips in or out of the field for any rotation angle choice, as it is shown in Section 3 for LOPS2. More in general, the range of galactic latitudes probed is always $0^{\circ}<|b|\lesssim 50^{\circ}$ regardless of $\varphi$, and therefore any property correlated with $b$, such as the amount of stellar crowding and the diversity of stellar populations probed (in terms of age, kinematics, metallicity, etc.) is approximately the same.
@@
Finally, it makes sense to investigate whether there is any way to rotate the field in order to move any extremely bright star ($V<1$) off the silicon. Paper I identified Deneb and Vega as the brightest sources on LOPN1, and Sirius and Canopus for LOPS1. Moving from LOPS1 to LOPS2 also implies that Sirius is now left outside the focal plane for any choice of $\varphi$, while there is no consequence on the other three stars, that are unavoidable and always monitored with the same number of NCAMs (18-24 for Vega, 6 for Deneb, 24 for Canopus) for almost any possible rotation angle. [^6] It could be possible in principle to fit Deneb within a $\sim 30^{\prime}$ CCD gap of the northern field at $\varphi\simeq 11^{\circ}$, but the technical feasibility of such a choice and its possible unintended consequences (see for instance Porterfield et al. 2016) need to be assessed. This will be done at a further stage, should LOPN1 be scheduled for observation. Fig. 15 (but also Section 3.2 below) investigates this point a bit further, looking at how the number of bright stars (up to $V<3$) in LOPS2 changes as a function of $\varphi$, at $10^{\circ}$ intervals from $0^{\circ}$ to $80^{\circ}$ (the symmetry of the PLATO field makes it invariant after $90^{\circ}$ rotations). The nominal $\varphi=0^{\circ}$ case is the only one for which all the five $21<V<$ 2 stars ($\delta$ CMa, $\epsilon$ CMa, $\gamma_{2}$ Vel, $\delta$ Vel, $\epsilon$ Car) always land on 6 NCAMs regions, while for instance four of them fall on 12 NCAMs in the $\varphi=40$-$50^{\circ}$ case. Finally, also the distribution of $32<V<$ 3 stars does not change significantly as a function of $\varphi$.
@@
In summary, we conclude that there is no reason to change the rotation angle of both LOPN1 and LOPS2 from its initial value of $\varphi=0^{\circ}$. We implicitly assume zero rotation (i.e., one side of the field almost tangent to the galactic plane) hereafter, as plotted in the all-sky projection in Fig. 2.
@@
Figure 2: All-sky Aitoff projection in galactic coordinates of LOPS2 and LOPN1, showing the formal constraints for the selection of the PLATO LOP fields and the synergies with other missions. The two pink circles represent the $|\beta|>63^{\circ}$ technical requirement for the center of the LOP fields (“allowed region”), implying that the overall envelopes of every allowed field choice are two ecliptic caps at $|\beta|\gtrsim 38^{\circ}$ (green circles). LOPS2 (lower left) and LOPN1 (upper right) are plotted with blue shades according to the number of co-pointing cameras, as in Fig. 1. The footprints of CoRoT (red), Kepler (magenta), and K2 (green) are over-plotted together with the TESS continuous viewing zone at $|\beta|\gtrsim 78^{\circ}$ (yellow circle). The background gray layer is color coded according to the areal density of $G<13.5$ stars from Gaia DR3. The celestial equator and poles are marked with a red line and crosses, respectively. This sky chart is also plotted in equatorial and Ecliptic coordinates in Fig. 17-18.
@@ 2.3 Synergies and follow-up
As for the previous space-based transit surveys, we first remind that no overlap between a LOP field and K2 or CoRoT (Howell et al., 2014; Auvergne et al., 2009) is possible at all, since the latter fields are all located close to the ecliptic by design, inaccessible during the LOP phase (Fig. 2). The TESS continuous viewing zone (CVZ) at $|\beta|>78^{\circ}$, on the other hand, is monitored at 100% in LOPN1 and $\gtrsim 90\%$ in LOPS2, mostly in six- and 12-camera regions. Finally, the Kepler field is fully included in the LOPN1, almost entirely with at least 12 NCAMs and up to 24. While the scientific case of following-up the Kepler TTV systems with PLATO is very solid (Jontof-Hutter et al., 2021a, b), we emphasize that it should not be seen as a main driver of the mission, but rather as a potential source of interesting additional science. [^7] It is worth mentioning that the follow-up of the Kepler field is part of the science goals of the proposed Earth-2.0 mission (Ge et al., 2022). Indeed, the last word is on the follow-up availability and timeliness, since the confirmation of the PLATO candidates and eventual mass determination is part of the mission products (and goals) through the Ground-based Observation Program (GOP) of the PLATO Consortium.
@@
It is worth mentioning two other important survey projects in the south that could have interesting implications for the PLATO preparation and/or follow-up: LSST (Ivezić et al., 2019) at the Vera C. Rubin Observatory, which will map the whole southern hemisphere (including LOPS2) beginning in 2025, and 4MOST at the VISTA telescope (de Jong et al., 2012), starting its consortium surveys also in 2024. The southern hemisphere will also host the Giant Magellan Telescope (GMT; Johns et al. 2012), which is scheduled to be operational in the early 2030s.
@@ 2.4 LOPS2 as first LOP field
 [^8] The rows give the coordinates for the field center in equatorial, galactic, and ecliptic reference frames and the number of targets available in the P1-P2-P4-P5 samples. The latter are calculated from PIC 2.0.0 assuming the “EOL 22” scenario (see Section 2.4).
Figure 3: LOPS2 footprint. Left panel: Orthographic projection in galactic coordinates centered on LOPS2. The field is color coded according to the number of NCAMs as in Fig. 1. The yellow circle at $\beta=-78^{\circ}$ approximates the southern TESS continuous viewing zone (CVZ), while the magenta circle is centered on the Large Magellanic cloud. Right panel: Same as in left panel but in equatorial coordinates.
@@
Figure 4: Interstellar extinction properties of LOPS2. Left panel: Average $V$-band extinction $A_{V}$ of the P1+P2 targets from PIC 2.0.0, calculated over HEALPix level-5 cells and represented in a color scale. Right panel: Same but for the P5 sample and over a finer level-6 grid. The high-extinction bump at $A_{V}\simeq 0.3$ is due to the closest parts of the Vela-Puppis region (Cantat-Gaudin et al., 2019).